OpenStudy (anonymous):

Use substitution to solve the system of equations. 5x + 5y = -30 2x - y = 15 (-2, 5) (12, -5) (3, -9) (-2, 10)

OpenStudy (anonymous):

<3 every one is looking at this question

OpenStudy (anonymous):

all 3 of us... :)

OpenStudy (anonymous):

there was like 5 of my friends looking now only 2 :(

OpenStudy (anonymous):

Isolate one of the variables in an equation

OpenStudy (anonymous):

what are the numbers in the () for?

OpenStudy (anonymous):

think those are answer choices

OpenStudy (anonymous):

Let me post ss

OpenStudy (anonymous):

Yes they are gand

OpenStudy (anonymous):

In the 2nd x=15+y/2 Take that x to the 1st

OpenStudy (anonymous):

3,-9

OpenStudy (anonymous):

So... first, isolate a variable. lets do x caus x is fun 5x+5y=-30 5x=-30-5y x=-6-y

OpenStudy (anonymous):

(3, -9)

OpenStudy (anonymous):

now, wherever u c an x in the second equation, plug in -6-y solve for y

OpenStudy (anonymous):

2(-6-y)-y=15 -12-2y-y=15 -12-3y=15 -3y=27 y=-9

OpenStudy (anonymous):

Now plug that into an equation and solve for x

OpenStudy (anonymous):

:O you lost me

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

idk where you got 2(-6-y)-y=15 -12-2y-y=15

OpenStudy (anonymous):

whats 2*-6 and 2*-y

OpenStudy (anonymous):

it's and x not *

OpenStudy (anonymous):

okay... you see where I got the first equation equal to x? x=-6-y -6-y is x so wherever u see an x, u can put in -6-y, because they're the same thing

OpenStudy (anonymous):

got it or still confused or did I do the wrong part...

OpenStudy (anonymous):

i think i got it but i don't get how to go from there :'(

OpenStudy (anonymous):

okay. 2x-y=15 2(-6-y)-y=15

OpenStudy (anonymous):

distribute the 2

OpenStudy (anonymous):

(3,-9)

OpenStudy (anonymous):

-12-2y-y=15

OpenStudy (anonymous):

with me so far?

OpenStudy (anonymous):

i kinda am lost on the steps but i kinda get what your doing

OpenStudy (anonymous):

K. want me to back up and show you again or keep going?

OpenStudy (anonymous):

:S idk brb i have to help my dad -_-

OpenStudy (anonymous):

k, c u l8r